The most recent financial statements for Crosby, Inc., follow. Sales for 2018 are projected to grow by 30 percent. Interest expense will remain constant; the tax rate and the dividend payout rate will also remain constant. Costs, other expenses, current assets, fixed assets, and accounts payable increase spontaneously with sales.

CROSBY, INC.
2017 Income Statement
  Sales $ 749,000
  Costs 584,000
  Other expenses 20,000
  Earnings before interest and taxes $ 145,000
  Interest paid 16,000
  Taxable income $ 129,000
  Taxes (21%) 27,090
  Net income $ 101,910
  Dividends $ 31,592
  Addition to retained earnings 70,318
CROSBY, INC.
Balance Sheet as of December 31, 2017
Assets Liabilities and Owners’ Equity
  Current assets   Current liabilities
    Cash $ 20,840     Accounts payable $ 55,000
    Accounts receivable 43,780     Notes payable 14,200
    Inventory 93,960       Total $ 69,200
      Total $ 158,580   Long-term debt $ 132,000
  Fixed assets   Owners’ equity
    Net plant and equipment $ 425,000     Common stock and paid-in surplus $ 115,500
    Retained earnings 266,880
      Total $ 382,380
  Total assets $ 583,580   Total liabilities and owners’ equity $ 583,580

If the firm is operating at full capacity and no new debt or equity is issued, what external financing is needed to support the 30 percent growth rate in sales?
Answer #1
CROSBY, INC.
2018 Income Statement
Sales [749000(1+.30)] 973700
cost    [584000(1+.30)] -759200
Other expense [20000(1+.30)] -26000
  Earnings before interest and taxes 188500
less:Interest -16000
Taxable income 172500
Tax [172500*.21] -36225
net income 136275
Dividend     [136275*31%] 42245
Addition to retained earning 94030

**Dividend payout ratio for 2017= 31592/101910= 31% rounded

Step 2)

Total asset for 2018 =Total asset for 2017 (1+% increase in sales)

              = 583580(1+.30)

               = 758654

step 3)Liabilities for 2018

Accounts payable [55000(1+.30)] 71500
Notes payable 14200
Long-term debt 132000
Total liabilities 217700

Step 4)

Owners equity for 2018 without EFN = common stock+Retained earning

       =115500 + (266880+94030)

       = 115500 + 360910

        = 476410

STEP 5)

External financing needed =Total asset -total liabilities -total equity

       = 758654-217700-476410

        = 64544